1a.If a product can be sold for $2,000 per unit and the variable cost per unit is $650, and fixed costs are $2 million, how many units must be produced and sold before the firm can break even? 1b. If an employee has a salary of $75,000 per year, benefits cost the company $7,200 per employee per year, and the employer’s share of the Social Security tax is 7.65 percent, what is the direct cost per year for this employee?

1c. A salesman in a recurring revenue firm is paid the equivalent of 2 months’ sales revenue for each new customer added. The charge for the service is $120 per month, and providing the service costs the company $50 per month per customer. It costs $25 to initially hook up each new customer. What would be the effect on this month’s expenses if the salesperson added fifty-five new customers this month?

1d. An assembly line can produce 125 units per hour and the material cost is $15 per unit. A single employee, who is paid $22 per hour, can operate the assembly line. The employee works an eight-hour day and there are twenty workdays in a month. The company’s contribution to Social Security is 7.65 percent of employee pay, and benefits cost the company about $565 per month per employee. Maintenance on the assembly line costs $1,000 per month. What is the total cost of producing the product for one month?
Answer #1

1a. Break Even Point = Fixed Cost / Contribution per unit

Contribution per unit = Sales Price - Variable Cost = $2,000 - $650 = $ 1350

Break Even Point = $ 2,000,000 / $1,350 = 1481.48 ie 1482 Units

Break Even Point 1,482 Units

1b.

Salary = $75,000

Add: Benefits provided = $7,200

Add: Social security tax (75,000 * 7.65%) = $5,737.5

Direct cost per year for an employee = $87,937.5

1c.

Amount paid to salesperson per customer = 2*120 = $240

Add Service cost = $50

Add Acquisition cost = $25

Total incremental cost for a customer acquired this month = $315

Total customer acquired this month = 55

Total incremental expenses = $17,325 ($ 315 * 55)

The expenses of firm will increase by $17,325 by adding 55 customers.

1d.

Total Number of working Days in a month

20

Total hours worked in a day

8

total hours in month

160

No of units produced in 1 Hour

125

No of units produced in a month

20000

Total Cost

Amount

Material cost

$        300,000

(15*20000)

Benefits cost

$                565

Maintenance on the assembly line costs

$            1,000

Employee Cost

$            3,520

(22*160)

Contribution to social security cost

$          269.28

(3520*7.65%)

Total Cost

$        305,354